# A Comprehensive Evaluation of Cognitive Biases in LLMs

### v1

@@ Abstract
We present a large-scale evaluation of 30 cognitive biases in 20 state-of-the-art large language models (LLMs) under various decision-making scenarios. Our contributions include a novel general-purpose test framework for reliable and large-scale generation of tests for LLMs, a benchmark dataset with 30,000 tests for detecting cognitive biases in LLMs, and a comprehensive assessment of the biases found in the 20 evaluated LLMs. Our work confirms and broadens previous findings suggesting the presence of cognitive biases in LLMs by reporting evidence of all 30 tested biases in at least some of the 20 LLMs. We publish our framework code to encourage future research on biases in LLMs: https://github.com/simonmalberg/cognitive-biases-in-llms.
@@ 1 Introduction
Transformer-based LLMs Vaswani (2017) and other Foundation Models (e.g., Gu and Dao, 2023) have gained significant attention in recent years. At an accelerating pace, models are becoming larger and more capable, conquering additional modalities such as vision and speech Shahriar et al. (2024). This makes LLMs increasingly attractive for complex reasoning Dziri et al. (2024); Saparov and He (2022) and decision-making tasks Eigner and Händler (2024); Echterhoff et al. (2024). However, using LLMs for high-stakes decision-making comes with severe risks, as they may produce flawed yet convincingly articulated outputs, such as hallucinations Zhang et al. (2023).
Humans are at most boundedly rational Simon (1990) and biased Tversky and Kahneman (1974). LLMs are trained on human-created data and typically fine-tuned on human-defined instructions Ouyang et al. (2022) and through Reinforcement Learning from Human Feedback (RLHF) Bai et al. (2022). Therefore, it is likely that human biases also creep into LLMs through the training procedure. While gender, ethical, and political biases in LLMs have been extensively studied Wan et al. (2023); Kamruzzaman et al. (2023); Bowen III et al. (2024); Rozado (2024), cognitive biases (e.g., see Figure 1) distorting human judgment and decision-making away from rationality Haselton et al. (2015) – a detrimental influence in high-stakes decision settings – have only very recently seen attention from LLM researchers.
@@
    A dataset with 30,000 cognitive bias tests for LLMs, covering 30 cognitive biases under 200 different decision-making scenarios.
@@ 3 Test Framework
Situation: Suppose you are [[a/an]] [[type]] manager at [[organization]]. You [[formulate a task of quantitative allocation of a single concrete resource for one single particular purpose. Do not include any numbers.]]. Prompt: Which allocation level do you choose for this purpose? Answer options: Option $1$: 00% Option $2$: $10$% Option $11$: $100$% Situation: Suppose you are [[a/an]] [[type]] manager at [[organization]]. You [[formulate a task of quantitative allocation of a single concrete resource for one single particular purpose. Do not include any numbers.]]. Prompt: Do you intend to allocate more than {{anchor}}for this purpose? Which allocation level do you choose for this purpose? Answer options: Option $1$: 00% Option $2$: $10$% Option $11$: $100$%
@@
We introduce a novel framework for reliably generating diverse and large-scale sets of tests for evaluating LLMs. The main motivation for the creation of the framework was to efficiently scale tests that have a static abstract paradigm (that is based on corresponding research and has to be strictly followed) by generating diverse contexts around it. The framework comprises four entities and three functions. Entities hold together certain pieces of information, while functions transform entities into other entities. All entities and functions are explained in the following. We use lower case letters $t,s,c,r,b,...$ to denote entities and their contents. Functions are denoted by upper case letters $G,D,E$. Some functions use an LLM internally. We use $f_{\theta}$ (inside $G$) or $h_{\theta}$ (inside $D$) to denote a pre-trained LLMs with parameters $\theta$.
@@ 3.1 Entities
A test case $c=[t_{1},t_{2},v,m]$ binds together two templates $t_{1}$ and $t_{2}$, a set of custom value generators $v=\{v_{1},v_{2},...\}$ and a metric $m$. $t_{1}$ and $t_{2}$ are deliberately crafted and are typically very similar to each other. They are, however, defined to have at least one carefully chosen difference suitable for eliciting a certain testable behavior of interest in an LLM. Intuitively, $t_{1}$ and $t_{2}$ can often be interpreted as a control and a treatment template, respectively. Custom value generators $v_{i}$ can be used to sample different values $w\sim\ v_{i}$ according to a specified distribution. Sampled custom values can then be inserted into template gaps, $x_{i}\leftarrow w,x_{i}\in g$. The metric $m$ defines the main estimation measure of the test outcome. A detailed description of a metric follows in Section 4.5. We denote test cases as $c^{\prime}$ when they include template instances $t_{1}^{\prime},t_{2}^{\prime}$ without any remaining gaps instead of raw templates $t_{1},t_{2}$.
@@ 3.2 Functions
A test generator $G(f_{\theta},c,s)$ takes an LLM $f_{\theta}$, a test case $c$, and a scenario $s$ to sample a test case $c^{\prime}\sim\ G(f_{\theta},c,s)$ by inserting values into the template gaps. These insertions can be either sampled from the LLM $f_{\theta}$ or from the custom value generators $\{v_{1},v_{2},...\}\in c$ according to the template instructions $p$ and scenario $s$. Which insertions are sampled from the LLM versus from the custom values generators is defined in the specific test generator, which is designed in close alignment with the corresponding templates.
In our framework implementation, the two template instances are sampled in two independent LLM calls $t_{1}^{\prime}\sim\ f_{\theta}^{GEN}(t_{1},s)$ and $t_{2}^{\prime}\sim\ f_{\theta}^{GEN}(t_{2},s)$, where $GEN$ denotes the particular LLM prompt used for generation (see Appendix D). However, identical gaps that exist in both templates, i.e., $g_{1}\cap g_{2},g_{1}\in t_{1},g_{2}\in t_{2}$, will only be filled once for $t_{1}$ and their insertions will then be copied over to $t_{2}$ to ensure consistency between the template instances. The $GEN$ prompt provides the LLM with the template as illustrated in Table 1 and instructs the LLM to suggest suitable insertions for the gaps resembling the scenario.
@@ 4.1 Bias Selection
Based on the available scientific literature, we designed a unique test case $c$ and corresponding test generator $G$ for each of the top 30 cognitive biases. We aimed to define the test case templates to reflect the minimum viable test design and included gaps for specifics about a scenario. An example test can be seen in Table 1. A detailed collection of scientific references and description of the exact test designs for all 30 biases can be found in Appendix B.
@@ 4.4 Dataset Validation
Secondly, we used the IFEval framework Zhou et al. (2023) to evaluate the instruction-following performance w.r.t. verifiable instructions (e.g., “Do not include any numbers.”). Test cases of 7 biases include instructions $p_{i}$ that contain constraints crucial for the cognitive biases’ testing designs, and IFEval thus allows us to fully validate the insertions of the respective gaps $x_{i}$ that the correctness of the corresponding tests is most dependent on. Among these 7 biases with verifiable instructions, 4 biases were generated 100% correctly, the other 3 biases’ populations have accuracies of 96.7%, 98.4%, and 99.6%. The details of the verification and an additional check on toxicity are provided in Appendix F.
@@ 5 Results & Discussion
Figure 5: The plot shows the absolute biasedness of models in relation to their size (bubble diameter) and Chatbot Arena score (as a measure of general capability). When no such score was available, we take the mean of the other models’ scores and mark the model with a ’*’.
A perspective on the absolute biasedness of the models in relation to other model characteristics such as size and general capability is provided in Figure 5. As a proxy for a model’s general capability, we show each model’s Chatbot Arena [^3] Chatbot Arena (scores from October 14, 2024) score on the horizontal axis. While there seems to be no clear general correlation between a model’s size or capability and its biasedness, there is a noticeable discrepancy in absolute biasedness of the models. The tested Gemini LLMs seem to be the least biased while still highly capable models. Qwen2.5 72B, GPT-4o mini, and Mistral Large follow up closely. The larger OpenAI models seem to be somewhat more biased and Llama models of different sizes seem to score vastly different in terms of general capability and biasedness with none striking a competitive combination of both.
@@
Figure 6 highlights clusters of models that exhibit similar biases. Some models that come from the same model families (e.g., Gemma, WizardLM) and some models of comparable size (e.g., Llama 3.2 1B and 3B) show similar bias characteristics. Further, four of the largest models tested can be found in the bottom four branches of the dendrogram, apparently showing similar behaviors.
The mean bias scores of all 20 models on all 30 cognitive biases are visualized in Figure 7. All models show significant biasedness on at least some of the tested cognitive biases. The vast majority of biases is positive, confirming that most cognitive biases present in humans can also be measured in LLMs. Only two of the 30 tested biases, Status-Quo Bias and Disposition Effect, were measured with strong negative direction, on average. On both biases, negative scores express a model’s preference for change. The Random models shows no biasedness on average, highlighting our metric’s strength as an unbiased estimator. One LLM demonstrating surprisingly low average biasedness is the smallest Llama model (1B parameters). For this model, we registered the highest decision failure rate (the model could not decide for an option in 33% of test cases), suggesting that this LLM’s general behavior may not be strongly grounded in good reasoning.
@@ 7 Limitations
Our paper provides a systematic framework for defining and conducting cognitive bias tests with LLMs. While we have demonstrated our pipeline using management decision-making as an example and established a respective dataset with 30,000 test cases for cognitive biases, our framework is theoretically generalizable beyond just this domain and task. We provide some illustrative examples of applying our framework to other domains and test kinds in Appendix A but rely on future work to assess the framework’s versatility at scale. Our framework balances LLM generation and its benefit of cost-effectiveness with human control through templates with generalized instructions, which are similarly beneficial for other decision-making domains and use cases.
@@ B.4 Social Desirability Bias
We obtain answers on a 7-point Likert scale $\sigma_{1}$. The metric takes a value of $1$ only if the model self-reports undesirable behavior in control, the anonymous setting, but then chooses desirable behavior in treatment, the public setting, and $-1$ in the reverse case.
@@ B.22 Endowment Effect
[…]
@@ B.23 Framing Effect
Our testing strategy follows directly from the attribute framing effect definition and replicates the study conducted in Fuenzalida et al. (2021). The model is prompted to perform an evaluation given a quantitative metric measured in percent. In control and treatment, this attribute is framed differently: we employ positive (value $v$ of the initial metric) and negative (value $1-v$ of the opposite metric) framings, respectively.
@@ Appendix F Analysis of the Dataset
Figure 9: Maximum Detoxify Hanu and Unitary team (2020) scores (those $>0.002$) reported for tests in our dataset. The highest toxicity score is obtained for Stereotyping, which is less than $0.02$. As the maximum Detoxify score is $1$, this result suggests that the contents of the dataset are largely non-toxic.

### v2

@@ Abstract
We present a large-scale evaluation of 30 cognitive biases in 20 state-of-the-art large language models (LLMs) under various decision-making scenarios. Our contributions include a novel general-purpose test framework for reliable and large-scale generation of tests for LLMs, a benchmark dataset with 30,000 tests for detecting cognitive biases in LLMs, and a comprehensive assessment of the biases found in the 20 evaluated LLMs. Our work confirms and broadens previous findings suggesting the presence of cognitive biases in LLMs by reporting evidence of all 30 tested biases in at least some of the 20 LLMs. We publish our framework code and dataset to encourage future research on cognitive biases in LLMs: https://github.com/simonmalberg/cognitive-biases-in-llms.
@@ 1 Introduction
Transformer-based LLMs Vaswani (2017) and other foundation models (e.g., Gu and Dao, 2023) have gained significant attention in recent years. At an accelerating pace, models are becoming larger and more capable, conquering additional modalities such as vision and speech Shahriar et al. (2024). This makes LLMs increasingly attractive for complex reasoning Dziri et al. (2024); Saparov and He (2022) and decision-making tasks Eigner and Händler (2024); Echterhoff et al. (2024). However, using LLMs for high-stakes decision-making, e.g., for managerial or public policy decisions, comes with severe risks, as they may produce flawed yet convincingly articulated outputs, such as hallucinations Zhang et al. (2023).
Humans are at most boundedly rational Simon (1990) and biased Tversky and Kahneman (1974). LLMs are trained on human-created data and typically fine-tuned on human-defined instructions Ouyang et al. (2022) and through reinforcement learning from human feedback (RLHF) Bai et al. (2022). Therefore, it is likely that human biases also creep into LLMs through the training procedure and data Caliskan et al. (2017). While gender, ethical, and political biases in LLMs have been extensively studied Wan et al. (2023); Kamruzzaman et al. (2023); Bowen III et al. (2024); Rozado (2024), cognitive biases distorting human judgment and decision-making away from rationality Haselton et al. (2015) have only very recently seen attention from LLM researchers. An example for a cognitive bias called the Framing Effect is illustrated in Figure 1. Interested readers will find detailed descriptions of 30 different cognitive biases in Appendix B. Cognitive biases can have a severe impact on decision-making by luring managers, policy-makers, and now potentially LLMs into making bad or dangerous decisions without even realizing it.
@@
    A dataset with 30,000 cognitive bias tests for LLMs, covering 30 cognitive biases under 200 different managerial decision-making scenarios.
@@ 3 Test Framework
Situation: Suppose you are [[a/an]] [[type]] manager at [[organization]]. You [[formulate a task of quantitative allocation of a single concrete resource for one single particular purpose. Do not include any numbers.]]. Prompt: Which allocation level do you choose for this purpose? Answer options: Option $1$: 0% Option $2$: 1010% Option $11$: 100100% Situation: Suppose you are [[a/an]] [[type]] manager at [[organization]]. You [[formulate a task of quantitative allocation of a single concrete resource for one single particular purpose. Do not include any numbers.]]. Prompt: Do you intend to allocate more than {{anchor}}% for this purpose? Which allocation level do you choose for this purpose? Answer options: Option $1$: 0% Option $2$: 1010% Option $11$: 100100%
@@
We introduce a novel framework for reliably generating diverse and large-scale sets of tests for evaluating LLMs. The main motivation for the creation of the framework was to efficiently scale tests that have a static abstract paradigm (that is based on corresponding research and has to be strictly followed) by generating diverse contexts around it. The framework comprises four entities and three functions. Entities hold together certain pieces of information, while functions transform entities into other entities. All entities and functions are explained in the following. We use lower case letters $t,s,c,r,b,$... to denote entities and their contents. Functions are denoted by upper case letters $G,D,E$. Some functions use an LLM internally. We use $f_{\theta}$ or $h_{\theta}$ to denote a pre-trained LLMs with parameters $\theta$.
@@ 3.1 Entities
A test case $c=[t_{1},t_{2},v,m]$ binds together two templates $t_{1}$ and $t_{2}$, a set of custom value generators $v=\{v_{1},v_{2},$...\} and a metric $m$. $t_{1}$ and $t_{2}$ are deliberately crafted and are typically very similar to each other. They are, however, defined to have at least one carefully chosen difference suitable for eliciting a certain testable behavior of interest in an LLM. Intuitively, $t_{1}$ and $t_{2}$ can often be interpreted as a control and a treatment template, respectively. Custom value generators $v_{i}$ can be used to sample different values $w\sim\ v_{i}$ according to a specified distribution. Sampled custom values can then be inserted into template gaps, $x_{i}\leftarrow w,x_{i}\in g$. The metric $m$ defines the main estimation measure of the test outcome. A detailed description of a metric follows in Section 4.5. We denote test cases as $c^{\prime}$ when they include template instances $t_{1}^{\prime},t_{2}^{\prime}$ without any remaining gaps instead of raw templates $t_{1},t_{2}$, i.e., all gaps have been filled.
@@ 3.2 Functions
A test generator $G(f_{\theta},c,s)$ takes an LLM $f_{\theta}$, a test case $c$, and a scenario $s$ to sample a test case $c^{\prime}\sim\ G(f_{\theta},c,s)$ by inserting values into the template gaps. These insertions can be either sampled from the custom value generators $\{v_{1},v_{2},...\}\in c$ or from the LLM $f_{\theta}$ according to the template instructions $p$ and scenario $s$. Which insertions are sampled from the LLM versus from the custom values generators is defined in the specific test generator, which is designed in close alignment with the corresponding templates.
In our framework implementation, the two template instances are sampled in two independent LLM calls $t_{1}^{\prime}\sim\ f_{\theta}^{GEN}(t_{1},s)$ and $t_{2}^{\prime}\sim\ f_{\theta}^{GEN}(t_{2},s)$, where $GEN$ denotes the particular LLM prompt used for generation (see Appendix D). However, identical gaps that exist in both templates will only be filled once for $t_{1}$ and their insertions will then be copied over to $t_{2}$ to ensure consistency between the template instances. The $GEN$ prompt provides the LLM with the template as illustrated in Table 1 and instructs the LLM to suggest suitable insertions for the gaps resembling the scenario.
@@ 4.1 Bias Selection
Based on the available scientific literature, we designed a unique test case $c$ and corresponding test generator $G$ for each of the top 30 cognitive biases. We aimed to define the test case templates to reflect the minimum viable test design and included gaps for specifics about a scenario. To ensure a high validity of the test designs, we conducted multiple rounds of internal peer reviews and subsequent revisions for all 30 tests until all authors of this work agreed that the test design was valid. An example test can be seen in Table 1. A detailed collection of scientific references and descriptions of the exact test designs for all 30 biases can be found in Appendix B.
@@ 4.4 Dataset Validation
Secondly, we used the IFEval framework Zhou et al. (2023) to evaluate the instruction-following performance w.r.t. verifiable instructions (e.g., “Do not include any numbers.”). Test cases of 7 biases include instructions $p_{i}$ that contain constraints crucial for the cognitive biases’ testing designs, and IFEval thus allows us to fully validate the insertions of the respective gaps $x_{i}$ that the correctness of the corresponding tests is most dependent on. Among these 7 biases with verifiable instructions, the percentage of tests where insertions satisfied the corresponding instruction was 100% for 4 biases and 96.7%, 98.4%, and 99.6% for the other 3 biases. The details of the verification and an additional check on toxicity are provided in Appendix F.
@@ 5 Results & Discussion
Figure 5: The plot shows the absolute biasedness (i.e., the strength of the biasedness, independent of direction) of models in relation to their size (bubble diameter) and Chatbot Arena score (as a measure of general capability). When no such score is available, we take the mean of the other models’ scores and mark the model with a ’*’.
A perspective on the absolute biasedness [^3] Absolute bias scores remove any leading signs to measure only strength and not the direction of the bias of the models in relation to other model characteristics such as size and general capability is provided in Figure 5. As a proxy for a model’s general capability, we show each model’s Chatbot Arena [^4] Chatbot Arena (scores from October 14, 2024) score on the horizontal axis. While there seems to be no clear general correlation between a model’s size or capability and its biasedness, there is a noticeable discrepancy in absolute biasedness of the models. The Gemini 1.5 Pro LLM seem to be the least biased while still highly capable model. Qwen2.5 72B, GPT-4o mini, Mistral Large, and Gemini 1.5 Flash follow up closely. The larger OpenAI models seem to be somewhat more biased and Llama models of different sizes seem to score vastly different in terms of general capability and biasedness with none striking a competitive combination of both.
@@
Figure 6 highlights clusters of models that exhibit similar biases. Some models that come from the same model families (e.g., Gemma, WizardLM) and some models of comparable size (e.g., Llama 3.2 1B and 3B) show similar bias characteristics. Further, three of the largest models tested can be found in the bottom three branches of the dendrogram, apparently showing similar behaviors.
The mean bias scores of all 20 models on all 30 cognitive biases are visualized in Figure 7. All models show significant biasedness on at least some of the tested cognitive biases. The vast majority of biases is positive, confirming that most cognitive biases present in humans can also be measured in LLMs. Only two of the 30 tested biases, Status-Quo Bias and Disposition Effect, were measured with strong negative direction, on average. On both biases, negative scores express a model’s preference for change. The Random model shows no biasedness on average, highlighting our metric’s strength as an unbiased estimator. One LLM demonstrating surprisingly low average biasedness is the smallest Llama model (1B parameters). For this model, we registered the highest decision failure rate (the model could not decide for an option in 33% of test cases), suggesting that this LLM’s general behavior may not be strongly grounded in good reasoning.
@@ 7 Limitations
Our paper provides a systematic framework for defining and conducting cognitive bias tests with LLMs. While we have demonstrated our pipeline using managerial decision-making as an example and established a respective dataset with 30,000 test cases for cognitive biases, our framework is theoretically generalizable beyond just this domain and task. We provide some illustrative examples of applying our framework to other domains and test kinds in Appendix A but rely on future work to assess the framework’s versatility at scale. Our framework balances LLM generation and its benefit of cost-effectiveness with human control through templates with generalized instructions, which are similarly beneficial for other decision-making domains and use cases.
@@ B.4 Social Desirability Bias
We obtain answers on a 7-point Likert scale $\sigma_{1}$. The metric takes a value of 11 only if the model self-reports undesirable behavior in control, the anonymous setting, but then chooses desirable behavior in treatment, the public setting, and 1-1 in the reverse case.
@@ B.22 Endowment Effect
In our experiment, we follow the latter approach Kahneman et al. (1990). In control, the LLM is prompted to evaluate the minimum amount it is willing to accept $\left(\text{WTA}\right)$ to give up the asset it owns. Symmetrically, in the treatment variant, we estimate the model’s maximum willingness to pay $\left(\text{WTP}\right)$ to acquire the same asset, which, in this case, it does not possess initially.
@@ B.23 Framing Effect
Our testing strategy follows directly from the attribute framing effect definition and replicates the study conducted in Fuenzalida et al. (2021). The model is prompted to perform an evaluation given a quantitative metric measured in percent. In control and treatment, this attribute is framed differently: we employ positive (value $v$ of the initial metric) and negative (value 1v1-v of the opposite metric) framings, respectively.
@@ Appendix F Analysis of the Dataset
Figure 9: Maximum Detoxify Hanu and Unitary team (2020) scores (those >0.002>0.002) reported for tests in our dataset. The highest toxicity score is obtained for Stereotyping, which is less than $0.02$. As the maximum Detoxify score is 11, this result suggests that the contents of the dataset are largely non-toxic.